OpenStudy (anonymous):

help please. simplify by combining like terms. xy+2x+4y+8/xy+2x+6y+12

OpenStudy (anonymous):

Is this what you are trying to symplify:\[xy+2x+4y+\frac{ 8 }{ xy }+2x+6y+12\]

OpenStudy (anonymous):

\[\frac{ xy+2x+4y+8 }{ xy+2x+6y+12 }\]

OpenStudy (alekos):

dont see how this could be done

OpenStudy (anonymous):

the answer i received from this problem solver says it is \[1- \frac{ 2 }{ x+6 }\] but i do not know how that became the answer

OpenStudy (alekos):

Not possible to get that answer from the original expression. can you send a photo of the question?

OpenStudy (alekos):

my mistake, it can be done. wasnt looking at it properly top line becomes (x+4)(y+2) bottom line (x+6)(y+2) then the y+2 terms cancel out leaving us with (x+4)/(x+6) then we can go (x+6-2)/(x+6) becomes (x+6)/(x+6) -2/(x+6) and presto 1 - 2/(x+6) !!! the answer given!!!

OpenStudy (anonymous):

thank you very much! you are such a life saver lol

OpenStudy (alekos):

no problem, you're welcome